OpenStudy (anonymous):

5y-10+y=7y+18-5y. Show work please. Solve and check. Label any contrdictions or identities.

OpenStudy (anonymous):

all numbers with "y" move them to one side

OpenStudy (anonymous):

5y + y - 7y + 5y = 18 + 10

OpenStudy (anonymous):

4y = 28

OpenStudy (anonymous):

y = 28/4

OpenStudy (anonymous):

y =7

OpenStudy (anonymous):

Where you have 4y= 28 was that suppose to be a 4 or 5

OpenStudy (anonymous):

5+1-7 +5 = 4